What is the geometric mean of the numbers 5 and 3125?
Answer by sdmmadam@yahoo.com(530)

What is the geometric mean of the numbers 5 and 3125?
The geometric mean between two numbers a and b is sqrt of (ab)
Therefore 5X3125 = 5X5X625 = (5^2)X(25)^2= (5X25)^2 = (125)^2
The geometric mean of 5 and 3125 is therefore sqrt of [(125)^2] = 125
Answer: 125